# Topological zero modes and bounded modes at smooth domain walls: Exact solutions and dualities

### v2

@@ Introduction
where $\bm{\Psi}(x)^{\dagger}=(\Psi(x)^{\dagger},\Psi(x))$ is a spinor with $\Psi(x)$ a fermion field, $\sigma_{xyz}$ the Pauli matrices, $p=-\mathrm{i}\partial$ the momentum operator, $\eta\geq 0$, and with $m(x),v(x)\in\mathbb{R}$. This equation describes a Dirac fermion of mass $m(x)$ and speed of light $2v(x)$ if $\eta=0$. In condensed matter, this equation describes a topological insulator with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and spin-orbit coupling $2v(x)$, or a spinless topological superconductor with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and $p$-wave superconducting pairing $\Delta_{p}(x)=2v(x)$. Here, we only focus on the zero modes $E=0$ which are normalizable and satisfy Dirichlet boundary conditions on infinite or semi-infinite intervals. These modes are necessarily eigenstates of $\sigma_{x}$ in the form
$\bm{\psi}(x)\propto\begin{pmatrix}1\\ s\end{pmatrix}\varphi(x),\,\,\sigma_{x}\bm{\psi}(x)=s\bm{\psi}(x),$ (2)
@@
where we absorbed $\eta$ into the definition of the fields $v(x)\to v(x)/\eta$ and $m(x)\to m(x)/\eta$. We refer to the eigenvalue $s$ as the pseudospin. Notice that the equation is real with real solutions and consequently the zero modes are Majorana modes analogous to the Majorana solutions of the Dirac equations. Hence, the real eigenstates of the equation above are also eigenstates of the antiunitary complex conjugation operator $K$ and of the antiunitary operator $C=K\sigma_{x}$, which can be interpreted as the antiunitary time-reversal and charge conjugation (or particle-hole symmetry in condensed matter).
@@
In the case where $v(x)$ and $m(x)$ can be expanded in powers of $y(x)=\left(1+\tanh{(x/2w)}\right)/2$ up to the first and second order, respectively, i.e., as $v(x)=v_{0}+v_{1}y(x)$, and $m(x)=m_{0}+m_{1}y(x)+m_{2}y(x)^{2}$, where $v_{0}=v_{L}$, $v_{1}=v_{R}-v_{L}$, $m_{0}=m_{L}$, $m_{1}+m_{2}=m_{R}-m_{L}$, and $m_{2}=2(m_{L}+m_{R})-4m_{D}$, with $m_{D}=m(0)$, we find an explicit general solution in terms of hypergeometric functions. For each set of exponents $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{\pm},\alpha_{R}^{\pm})$ the equation becomes
@@
1) general case $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left({\alpha_{L}^{\pm}}-{\alpha_{R}^{\mp}}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{w\left(\alpha_{L}^{\pm}+\alpha_{R}^{\mp}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$=\dfrac{\mathrm{e}^{\left[\pm\left(q_{L}+q_{R}\right)-s\left(v_{L}+v_{R}\right)\right]\,x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{w\left[\pm\left(q_{L}-{q_{R}}\right)+sv_{1}\right]}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a_{1,2}=\pm w\left({q_{L}}-{q_{R}}\right)+\frac{1}{2}+\frac{1}{2}\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}$
$b_{1,2}=\pm w\left({q_{L}}-{q_{R}}\right)+\frac{1}{2}-\frac{1}{2}\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}$
$c_{1,2}=1+2wq_{L,R}$
2) $v(x)=v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{-sv\,x}\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}+\frac{1}{2}\sqrt{1+4w^{2}m_{2}}$
$b_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}-\frac{1}{2}\sqrt{1+4w^{2}m_{2}}$
$c_{1,2}=1+2wq_{L,R}$
3) $v(x)=v$, $m_{L,R}=m$ $\varphi_{1,2}^{s}(x)=\mathrm{e}^{\left(-sv\pm\sqrt{{v}^{2}+m}\right)\,x}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\sqrt{1+4w^{2}m_{2}}$,  $c=1+2{w\sqrt{{v}^{2}+m}}$
4) $v(x)=v$, $m(x)=m$ $\varphi_{1,2}^{s}(x)=\mathrm{e}^{\left(-sv\pm\sqrt{{v}^{2}+m}\right)\,x}$
5) $v_{L,R}=\pm v$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)-2wsv}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}+\frac{1}{2}\sqrt{\left(4wsv+1\right)^{2}+4w^{2}m_{2}}$
$b_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}-\frac{1}{2}\sqrt{\left(4wsv+1\right)^{2}+4w^{2}m_{2}}$
$c_{1,2}=1+2wq_{L,R}$
6) $v_{L,R}=\pm v$, $m_{L,R}=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\sqrt{\left(4wsv+1\right)^{2}+4w^{2}m_{2}}$,  $c=1+2{w\sqrt{{v}^{2}+m}}$
7) $v_{L,R}=\pm v$, $m(x)=m$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\pm\sqrt{{v}^{2}+m}\,\,x}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{-2wsv}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\left(4wsv+1\right)$,  $c=1+2{w\sqrt{{v}^{2}+m}}$
8) $q_{L,R}=q$ $\varphi_{1,2}^{s}(x)=\dfrac{\mathrm{e}^{\left[\pm 2q-s(v_{L}+v_{R})\right]\,x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{wsv_{1}}}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\sqrt{\left(2wsv_{1}-1\right)^{2}+4w^{2}m_{2}}$,   $c=1+2wq$,
@@
These topological invariants depend only on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$. The phase on the left $x\ll-w$ is topologically nontrivial if $m_{L}\leq 0$ and $v_{L}\neq 0$ with topological invariant equal to $W_{L}=\operatorname{sgn}(v_{L})$ and trivial otherwise. Analogously, the phase on the right $x\gg w$ is nontrivial if $m_{R}\leq 0$ and $v_{R}\neq 0$ with $W_{R}=\operatorname{sgn}(v_{R})$ and trivial otherwise (see Refs. 14, 8).
@@
infinite interval $-\infty<x<\infty$
$W_{L}W_{R}=-1$ 2 solutions $\begin{cases}\varphi_{1}^{s}(x),\\ \varphi_{2}^{s}(x),\end{cases}$ $s=-W_{L}=W_{R}$
$\operatorname{sgn}(m_{L})=\operatorname{sgn}(m_{R})=-1$ and $v_{L}v_{R}<0$
$W_{L}=0$, $|W_{R}|=1$
$\operatorname{sgn}(m_{L})=1$, $\operatorname{sgn}(m_{R})=-1$ with $v_{R}\neq 0$
1 solution $\varphi_{1}^{s}(x)$ $s=W_{R}$
$W_{L}W_{R}=-1$ (limiting case)
$\operatorname{sgn}(m_{L})=0$, $\operatorname{sgn}(m_{R})=-1$ with $v_{L}v_{R}<0$
$|W_{L}|=1$, $W_{R}=0$
$\operatorname{sgn}(m_{L})=-1$, $\operatorname{sgn}(m_{R})=1$ with $v_{L}\neq 0$
1 solution $\varphi_{2}^{s}(x)$ $s=-W_{L}$
$W_{L}W_{R}=-1$ (limiting case)
$\operatorname{sgn}(m_{L})=-1$, $\operatorname{sgn}(m_{R})=0$ with $v_{L}v_{R}<0$
otherwise 0 solutions
semi-infinite interval $0\leq x<\infty$
$|W_{R}|=1$ 1 solution $\sum_{i}A_{i}\varphi_{i}^{s}(x)$ $s=W_{R}$
$\operatorname{sgn}(m_{R})=-1$
otherwise 0 solutions
@@
Therefore, the existence and number of zero modes and their pseudospin in the infinite interval $-\infty<x<\infty$ only depend on the topological invariants $W_{L,R}$, i.e., on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$, with their pseudospin determined only by $W_{L,R}$. These modes are localized at the smooth domain wall, and their number coincides with the difference between the topological invariants on the left and right sides $|W_{R}-W_{L}|$, as expected from the bulk-boundary correspondence[11, 12, 13]. Analogously, the existence of a zero mode in the semi-infinite interval $0\leq x<\infty$ only depends on the topological invariant $W_{R}$, i.e., on the values of the fields $m_{R}$ and $v_{R}$ at $x>w$, with its pseudospin determined only by $W_{R}$. This mode is localized at an infinite wall at $x=0$, when the topological invariant on the right side is nontrivial $|W_{R}|=1$, as expected from the bulk-boundary correspondence.
@@
On the contrary, zero modes localized at a smooth domain wall have "hair" in the sense that they are nonfeatureless, i.e., cannot be described entirely by a small set of numbers, at least not at short length scales. Indeed, in the case $w<\xi,\lambda$, modes localized at a smooth domain wall appear featureless at large distances $|x|>\xi,\lambda>w$ but not at short distances $|x|<w$. As long as they are looked at from a large distance $|x|>\xi,\lambda>w$, these modes are effectively hairless, i.e., they can be treated as effectively featureless modes at long length scales, being as good as zero modes localized at sharp phase boundaries, and being characterized only by a few numbers ($\xi,\lambda$, or $\alpha$). We refer to this case as zero modes with "short hair". Dramatically different is the case where $w>\xi,\lambda$. In this case, the features of the wavefunctions cannot be neglected since the length scales associated with these features are larger than the decaying lengths of the modes. Hence, these wavefunctions are nonfeatureless at all length scales, i.e., cannot be described only by a set of few numbers $\xi,\lambda$, or $\alpha$ at any scale. We refer to this case as zero modes with "long hair". Another way to look at the difference between the case $w<\xi,\lambda$ and the case $w>\xi,\lambda$ is to say that the spatial extension of the mode is larger than the spatial extension of the smooth domain wall. Notice also that the property of being a short-hair $w<\xi,\lambda$ or long-hair $w>\xi,\lambda$ mode depends on whether we look at the mode from the left side (i.e., $x<0$) or from the right side (i.e., $x>0$ for a smooth domain wall at $x=0$). We notice that, even though their topological nature is not apparent at all length scales, the existence of the short-hair $w<\xi,\lambda$ and long-hair $w>\xi,\lambda$ modes at smooth domain walls is still a direct consequence of the bulk-boundary correspondence. In the context of one-dimensional topological superconductors, featureless, hairless zero modes correspond to textbook examples of Majorana zero modes localized at a sharp boundary between a topologically nontrivial phase and a trivial phase (e.g., the vacuum), whereas nonfeatureless zero modes in the short-hair or long-hair regimes correspond to different incarnations of quasi-Majorana modes which appear at smooth boundaries[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
1) general case $\widetilde{\varphi}_{1,2}(x)=\dfrac{\mathrm{e}^{\pm\left(q_{L}+q_{R}\right)x/2}}{\left(2\cosh{\left(\frac{x}{2w}\right)}\right)^{\pm w\left(q_{L}-q_{R}\right)}}{{\,}_{2}F_{1}\left(a_{1,2},b_{1,2},c_{1,2},\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}+\frac{1}{2}\sqrt{1+4w^{2}V_{2}}$
$b_{1,2}=\pm w({q_{L}}-{q_{R}})+\frac{1}{2}-\frac{1}{2}\sqrt{1+4w^{2}V_{2}}$
$c_{1,2}=1+2wq_{L,R}$
2) $V_{L,R}=V$ $\widetilde{\varphi}_{1,2}(x)=\mathrm{e}^{\pm\sqrt{V-E}\,\,x}{{\,}_{2}F_{1}\left(a,b,c,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\sqrt{1+4w^{2}V_{2}}$,  $c=1+2{wq}$,  $q=q_{L,R}$
3) $V(x)=V$ $\widetilde{\varphi}_{1,2}(x)=\mathrm{e}^{\pm\sqrt{V-E}\,\,x}$
4) $q_{L,R}=0$, $V_{L,R}=E$ $\widetilde{\varphi}_{1,2}(x)={{\,}_{2}F_{1}\left(a,b,1,\tfrac{1}{\mathrm{e}^{\mp\,x/w}+1}\right)}$
$a,b=\frac{1}{2}\pm\frac{1}{2}\sqrt{1+4w^{2}V_{2}}$
@@
infinite interval $-\infty<x<\infty$
$\widetilde{W}_{L}=\widetilde{W}_{R}=1$ 2 solutions $\begin{cases}\widetilde{\varphi}_{1}(x)\\ \widetilde{\varphi}_{2}(x)\end{cases}$
$E\geq V_{L},V_{R}$
$\widetilde{W}_{L}=0$, $\widetilde{W}_{R}=1$ 1 solution $\widetilde{\varphi}_{1}(x)$
$V_{L}>E\geq V_{R}$
$\widetilde{W}_{L}=1$, $\widetilde{W}_{R}=0$ 1 solution $\widetilde{\varphi}_{2}(x)$
$V_{R}>E\geq V_{L}$
otherwise 0 solutions
semi-infinite interval $0\leq x<\infty$
$\widetilde{W}_{R}=1$ 1 solution $\sum_{i}A_{i}\widetilde{\varphi}_{i}(x)$
$E\geq V_{R}$
otherwise 0 solutions
@@
We consider the duality transformation in Eq. 15, giving $v(x)^{2}+sv^{\prime}(x)+m(x)=V(x)-E$. This transformation mandates that, if $\widetilde{W}_{L,R}=1$ then $|W_{L,R}|=1$ assuming $v_{L,R}\neq 0$, since in this case $v_{L,R}^{2}+m_{L,R}=V_{L,R}-E\leq 0$. The converse is not always true, since $|W_{L,R}|=1$ mandates $\widetilde{W}_{L,R}=1$ only if $|m_{L,R}|>v_{L,R}^{2}$. Hence, the 2 bounded modes $\widetilde{\varphi}_{1,2}(x)$ on the infinite interval $-\infty<x<\infty$ with $\widetilde{W}_{L}=\widetilde{W}_{R}=1$, i.e., with $E\geq V_{L,R}$ correspond via the duality transformation to the 2 zero modes $\varphi_{1,2}^{s}(x)$, if we assume $v_{L}v_{R}<0$, since in this case $|W_{L,R}|=1$ and $W_{L}W_{R}=-1$. The two bounded modes and the zero modes share the same oscillatory character, with the two bounded modes oscillating and the two zero modes decaying exponentially with damped oscillations and decay rates $|v_{L,R}|$. On each side of the barrier, these oscillations have the same momenta, which are given by $\kappa_{L,R}=\mathrm{Im}(\sqrt{v_{L,R}^{2}+m_{L,R}})=\mathrm{Im}(\sqrt{V_{L,R}-E})$. In the limiting cases where either of these momenta vanishes, the bounded modes converge to a constant value and the zero nodes decay exponentially without oscillations, respectively on the left (if $E=V_{L}$), on the right (if $E=V_{R}$), and on both sides (if $E=V_{L}=V_{R}$).
@@
infinite interval $-\infty<x<\infty$
$\widetilde{W}_{L}=\widetilde{W}_{R}=1$ $W_{L}W_{R}=-1$
$V_{L,R}\leq E$ $v_{L,R}^{2}+m_{L,R}\leq 0$
choosing $v_{L}v_{R}<0$
2 wavelike modes $\widetilde{\varphi}_{1,2}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 2 zero modes $\varphi_{1,2}^{s}$
$V_{L}=E\Rightarrow$ const. at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{L}^{2}+m_{L}=0\Rightarrow$ exp. decay at $\infty$- ∞
$V_{L}<E\Rightarrow$ oscill. at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{L}^{2}+m_{L}<0\Rightarrow$ exp. damped oscill. at $\infty$- ∞
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞
$\widetilde{W}_{L}=0$, $\widetilde{W}_{R}=1$ $W_{L}=0$, $|W_{R}|=1$
$V_{L}>E\geq V_{R}$ $v_{L}^{2}+m_{L}>0\geq v_{R}^{2}+m_{R}$
choosing $|m_{L}|>v_{L}^{2}$ and $v_{R}\neq 0$
1 Schokley-like mode $\widetilde{\varphi}_{1}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 1 zero mode $\varphi_{1}^{s}$
exp. decay at $\infty$- ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ exp. decay at $\infty$- ∞
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞
semi-infinite interval $0\leq x<\infty$
$\widetilde{W}_{R}=1$ $|W_{R}|=1$
$V_{R}\leq E$ $v_{R}^{2}+m_{R}\leq 0$
choosing $v_{R}\neq 0$
1 wavelike mode $\sum A_{i}\widetilde{\varphi}_{i}$ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ 1 zero mode $\sum A_{i}\varphi_{i}^{s}$
$V_{R}=E\Rightarrow$ const. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}=0\Rightarrow$ exp. decay at +$\infty$+ ∞
$V_{R}<E\Rightarrow$ oscill. at +$\infty$+ ∞ $\overleftrightarrow{\qquad\qquad\qquad\qquad\qquad\qquad\qquad}$ $v_{R}^{2}+m_{R}<0\Rightarrow$ exp. damped oscill. at +$\infty$+ ∞

### v4

@@ Introduction
where $\Psi(x)^{\dagger}=(\psi(x)^{\dagger},\psi(x))$ is a spinor, $\sigma_{xyz}$ the Pauli matrices, $p=-{\mathrm{i}}\partial$ the momentum operator, $\eta\geq 0$, and with $m(x),v(x)\in\mathbb{R}$. This equation describes a Dirac fermion of mass $m(x)$ and speed of light $2v(x)$ if $\eta=0$. In condensed matter, this equation describes a topological insulator with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and spin-orbit coupling $2v(x)$, or a spinless topological superconductor with effective mass $1/2\eta$, chemical potential $\mu(x)=-m(x)$, and $p$-wave superconducting pairing $\Delta_{p}(x)=2v(x)$. Here, we only focus on the zero modes $E=0$ which are normalizable and satisfy Dirichlet boundary conditions on infinite or semi-infinite intervals. These modes are necessarily eigenstates of $\sigma_{x}$ in the form
$\Psi(x)\propto\begin{pmatrix}1\\ s\end{pmatrix}\varphi(x),\quad\sigma_{x}\Psi(x)=s\Psi(x),$ (2)
@@
where we absorbed $\eta$ into the definition of the fields $v(x)\to v(x)/\eta$ and $m(x)\to m(x)/\eta$. We refer to the eigenvalue $s$ as the pseudospin. Notice that the equation is real with real solutions and consequently the zero modes are Majorana modes analogous to the Majorana solutions of the Dirac equations. Hence, the real eigenstates of the equation above are also eigenstates of the antiunitary complex conjugation operator $K$ and of the antiunitary operator $C=K\sigma_{x}$, which can be interpreted as the antiunitary time-reversal and charge conjugation (or particle-hole symmetry in condensed matter), respectively. Notice that Eq. 1 can be unitary transformed via $v(x)p\sigma_{y}\to v(x)p\left(\sin{\theta}\,\sigma_{y}+\cos{\theta}\,\sigma_{x}\right)$, with $\theta$ an arbitrary gauge angle. In the specific case of a superconductor, this corresponds to the gauge invariance with respect to the global phase of the superconductor. Moreover, one can consider a more general equation where the gauge angle is position dependent $\theta\to\theta(x)$ giving $v(x)p\sigma_{y}\to v(x)p\left(\sin{\theta(x)}\,\sigma_{y}+\cos{\theta(x)}\,\sigma_{x}\right)$. In this case, however, the solutions cannot be written in terms of eigenstates of a single Pauli matrix for all values of the position $x$ as in Eq. 2. We find numerical evidence that for $\theta^{\prime}(x)\neq 0$, there are no zero energy solutions of the Jackiw-Rebbi equation unless in the limiting case $m(x)=0$. This limiting case corresponds to the closing of the bulk gap of the topological insulator or topological superconductor, yielding a continuous spectra around zero energy, and no localized zero-modes.
@@
[…]
@@
These topological invariants depend only on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$. The phase on the left $x\ll-w$ is topologically nontrivial if $m_{L}\leq 0$ and $v_{L}\neq 0$ with topological invariant equal to $W_{L}=\operatorname{sgn}(v_{L})$ and trivial otherwise. Analogously, the phase on the right $x\gg w$ is nontrivial if $m_{R}\leq 0$ and $v_{R}\neq 0$ with $W_{R}=\operatorname{sgn}(v_{R})$ and trivial otherwise (see Refs. [14, 8]).
@@
Therefore, the existence and number of zero modes and their pseudospin in the infinite interval $-\infty<x<\infty$ only depend on the topological invariants $W_{L,R}$, i.e., on the values of the fields $m_{L,R}$ and $v_{L,R}$ at large distances $|x|>w$. These modes are localized at the smooth domain wall, and their number coincides with the difference between the topological invariants on the left and right sides $|W_{R}-W_{L}|$, as expected from the bulk-boundary correspondence [11, 12, 13]. Analogously, the existence of a zero mode in the semi-infinite interval $0\leq x<\infty$ only depends on the topological invariant $W_{R}$, i.e., on the values of the fields $m_{R}$ and $v_{R}$ at $x>w$, with its pseudospin determined only by $W_{R}$. This mode is localized at an infinite wall at $x=0$, when the topological invariant on the right side is nontrivial $|W_{R}|=1$, as expected from the bulk-boundary correspondence.
@@
On the contrary, zero modes localized at a smooth domain wall have "hair" in the sense that they are nonfeatureless, i.e., cannot be described entirely by a small set of numbers, at least not at short length scales. Indeed, in the case $w<\xi,\lambda$, modes localized at a smooth domain wall appear featureless at large distances $|x|>\xi,\lambda>w$ but not at short distances $|x|<w$. As long as they are looked at from a large distance $|x|>\xi,\lambda>w$, these modes are effectively hairless, i.e., they can be treated as effectively featureless modes at long length scales, being as good as zero modes localized at sharp phase boundaries, and being characterized only by a few numbers ($\xi,\lambda$, or $\alpha$). We refer to this case as zero modes with "short hair". Dramatically different is the case where $w>\xi,\lambda$. In this case, the features of the wavefunctions cannot be neglected since the length scales associated with these features are larger than the decaying lengths of the modes. Hence, these wavefunctions are nonfeatureless at all length scales, i.e., cannot be described only by a set of few numbers $\xi,\lambda$, or $\alpha$ at any scale. We refer to this case as zero modes with "long hair". Notice also that the property of being a short-hair $w<\xi,\lambda$ or long-hair $w>\xi,\lambda$ mode depends on whether we look at the mode from the left side (i.e., $x<0$) or from the right side (i.e., $x>0$ for a smooth domain wall at $x=0$). We notice that, even though their topological nature is not apparent at all length scales, the existence of the short-hair $w<\xi,\lambda$ and long-hair $w>\xi,\lambda$ modes at smooth domain walls is still a direct consequence of the bulk-boundary correspondence. In the context of one-dimensional topological superconductors, featureless, hairless zero modes correspond to textbook examples of Majorana zero modes localized at a sharp boundary between a topologically nontrivial phase and a trivial phase (e.g., the vacuum), whereas nonfeatureless zero modes in the short-hair or long-hair regimes correspond to different incarnations of quasi-Majorana modes which appear at smooth boundaries [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
We consider the duality transformation in Eq. 15, giving $v(x)^{2}+sv^{\prime}(x)+m(x)=V(x)-E$. This transformation mandates that, if $\widetilde{W}_{L,R}=1$ then $|W_{L,R}|=1$ assuming $v_{L,R}\neq 0$, since in this case $v_{L,R}^{2}+m_{L,R}=V_{L,R}-E\leq 0$. The converse is not always true, since $|W_{L,R}|=1$ mandates $\widetilde{W}_{L,R}=1$ only if $|m_{L,R}|>v_{L,R}^{2}$. Hence, the 2 bounded modes $\widetilde{\varphi}_{1,2}(x)$ on the infinite interval $-\infty<x<\infty$ with $\widetilde{W}_{L}=\widetilde{W}_{R}=1$, i.e., with $E\geq V_{L,R}$ correspond via the duality transformation to the 2 zero modes $\varphi_{1,2}^{s}(x)$, if we assume $v_{L}v_{R}<0$, since in this case $|W_{L,R}|=1$ and $W_{L}W_{R}=-1$. The two bounded modes and the zero modes share the same oscillatory character, with the two bounded modes oscillating and the two zero modes decaying exponentially with damped oscillations and decay rates $|v_{L,R}|$. On each side of the barrier, these oscillations have the same momenta, which are given by $\kappa_{L,R}=\mathrm{Im}(\sqrt{v_{L,R}^{2}+m_{L,R}})=\mathrm{Im}(\sqrt{V_{L,R}-E})$. In the limiting cases where either of these momenta vanishes, the bounded modes converge to a constant value and the zero modes decay exponentially without oscillations, respectively on the left (if $E=V_{L}$), on the right (if $E=V_{R}$), and on both sides (if $E=V_{L}=V_{R}$).